OpenStudy (anonymous):

(MEDAL & FAN) If 9(x - 9) = -11, then x = ? -92/9 70/9 -2/9 -20/9

OpenStudy (denonakavro):

9(x-9) = -11 9x-81 = -11 9x= 70 x=70/9 x=7.7778

OpenStudy (anonymous):

k

OpenStudy (crabbyoldgamer):

Or you can plug in the answer choices to see which one works.